# ProGeny II: the impact of libraries and model configurations on inferred galaxy properties in SED fitting

### v1

@@ Abstract
We use a volume-complete sample of $\sim$8,000 galaxies from the GAMA survey to characterise the impact of stellar population libraries (SPLs) and model configurations on the resulting inferred galaxy properties from Spectral Energy Distribution (SED) fitting. We compare a fiducial SPL from ProGeny (a new tool that can generate SPLs quickly and flexibly) against five other commonly used SPLs using the SED-fitting code ProSpect. The impact of selecting each SPL is compared to the consequence of changing the model implementation in the SED fitting process, including the implementation of metallicity evolution versus a fixed or constant metallicity, and a functional parametric star formation history (SFH) versus a stepwise parametric (or “non-parametric”) SFH. Furthermore, we use ProGeny to assess the impact of sub-SPL choices, including isochrone selection, atmosphere selection, and IMF selection. Through a comparison of derived stellar masses, star formation rates, metallicities, ages, and the inferred cosmic star formation history (CSFH), we rank the impact of varying choices. Overall the assumption of a solar metallicity creates the greatest biases, with a substantial impact also caused by the choice of a specific SPL. To recover a CSFH most consistent with observations, we advocate for the use of the fiducial implementation with a skewed Normal functional form for the SFH, and an evolving metallicity. While all current SPLs currently underestimate the peak in the CSFH, ProGeny and FSPS are the closest to recovering the observed CSFH.
@@ 1 Introduction
Stellar population sysnthesis (SPS) modelling uses knowledge about the evolutionary track of stars (isochrones), combined with a stellar initial mass function (IMF, for example Salpeter, 1955; Kroupa, 2002; Chabrier, 2003) and a set of stellar atmospheres that provide spectra at every evolutionary stage of a star to build up a stellar population generatively (for example, Tinsley & Gunn, 1976; Tinsley, 1978; Bruzual & Charlot, 1993).
SPS models are capable of producing what are known as Simple Stellar Populations (SSPs), which for a given age and metallicity provide a template for the stellar population. A library of such SSPs should cover the entire stellar popualtion, and is referred to as a Stellar Population Library (SPL). These SPLs (of which many flavours exist, including for example Leitherer et al., 1999; Bruzual & Charlot, 2003; Maraston, 2005; Vazdekis et al., 2016) are used as a foundational input to model the observed Spectral Energy Distributions (SEDs) of galaxies via their broadband photometry, in a process named SED fitting. In SED fitting, a stellar population is described in terms of a star formation history (SFH) to account for its age distribution, and a corresponding metallicity history (ZH). This description is used to recreate the stellar emission of the galaxy by building up the appropriate SSPs. An additional dust component is typically modelled, which facilitates the attenuation of the stellar light, as well as the re-emission of this light in the far-infrared through an energy balance process. Much has been written on the topic of SED fitting, and some excellent reviews on the topic have been provided by Walcher et al. (2011); Conroy (2013). Today, many different SED fitting codes exist in the literature, including Magphys (da Cunha et al., 2008), BayeSED (Han & Han, 2012), CIGALE (Boquien et al., 2019), Prospector (Leja et al., 2017), BAGPIPES (Carnall et al., 2018), BEAGLE (Chevallard & Charlot, 2016), ProSpect (Robotham et al., 2020).
@@
Comparisons of the galaxy properties to have been derived by different SED fitting codes have been made in the past (such as the extensive analysis by Pacifici et al., 2023), and furthermore many studies have analysed the impact of individual elements of SED fitting. This includes Carnall et al. (2019); Leja et al. (2019); Lower et al. (2020); Suess et al. (2022) who studied the impact of SFH parametrisations on stellar masses using Prospector and BAGPIPES, Thorne et al. (2022a) who studied the impact of the inclusion of AGN as a separate component, Haskell et al. (2024) who studied the impact of star formation burstiness, and Jones et al. (2022) who studied the impact of different dust models.
@@
In this work, we aim to characterise the impact of the selected SPL on the final galaxy properties from SED fitting. We compare the SPL impact to that of other modelling choices, including the SFH parametrisation, and the metallicity modelling implementation. Finally, by generating individual SPLs with typical stellar atmospheres, isochrones, and IMFs, we further characterise the impact of these sub-SPL choices on the final galaxy properties derived using SED fitting. This sub-SPL comparison is facilitated through the development of a new SPL generation software ProGeny, which is outlined in a companion paper by Robotham & Bellstedt (in prep).
@@ 2.1 GAMA
Vmax values for each galaxy in the sample have been computed using ProSpect, by which the maximum redshift is computed where the galaxy falls within the $m_{r}<19.65$ limit.
@@ 2.2 Stellar Population Libraries
SPL Reference Isochrone Atmosphere IMF Z Range
@@
M05 Maraston (2005) Maraston Kroupa (2002) [1e-3, 0.04]
FSPS Conroy & Wechsler (2009) Padova MILES Chabrier (2003) [2e-4, 0.03]
@@
ProGeny Robotham & Bellstedt (in prep) MIST C3K Chabrier (2003) [2e-6, 0.063]
  • Isochrones: Padova 1994 (Girardi et al., 2002), Geneva (Lejeune & Schaerer, 2001), Maraston (Maraston, 2005), Padova (Marigo & Girardi, 2007; Marigo et al., 2008), Cambridge STARS (Eldridge et al., 2008), PARSEC-COLIBRI (Marigo et al., 2017), MIST (Dotter, 2016).
  • Stellar atmospheres: BaSeL (Lejeune et al., 1997), STELIB (Le Borgne et al., 2003), Pickles (Pickles, 1998), MILES (Sánchez-Blázquez et al., 2006), C3K (Conroy et al., 2014).
The six SPLs compared in this work are presented in Table 1, and include (in order of original release) BC03 (Bruzual & Charlot, 2003), M05 (Maraston, 2005), FSPS (Conroy & Wechsler, 2009), BPASS (Stanway & Eldridge, 2018), CB19 (an update of Bruzual & Charlot 2003, and introduced in Plat et al. 2019), and ProGeny [^3] Available publicly at https://github.com/asgr/ProGeny (Robotham & Bellstedt in prep) [^4] While available for use within ProSpect, we have elected not to include the libraries EMILES (Vazdekis et al., 2016) and XSL (Verro et al., 2022) in this paper, due to their truncated wavelength range, and lack of availability of low-age templates. In general, these libraries are better suited for spectral analysis, rather than broadband SEDs. . Each SPL has been constructed using stellar atmospheres and isochrones that originate from slightly different sources. Furthermore, the nature by which these libraries have been constructed (including interpolation schemes, and how different atmosphere sets are prioritised) also vary. For further details on the exact construction of these SPLs, please see Robotham & Bellstedt (in prep).
@@ 3 Method
We use the SED fitting code ProSpect (Robotham et al., 2020) to test and compare the relative impact of using each of the independent stellar population libraries (SPLs herein), the metallicity implementation (as this is a parameter that varies significantly in terms of its degeneracy with age with different SPLs), and broadly the SFH parametrisation.
@@
The final b5 SFH configuration employs the same metallicity evolution implementation as the fiducial model, however it employs a stepwise parametrisation for the star formation history (available in ProSpect as the massfunc_b5 function), more similar to the stepwise parametric SFHs used typically by codes like Prospector (Leja et al., 2017). This parametriation divides the SFH into five time bins, where the star formation rate (SFR) is fitted independently in each bin. The advantage of this over massfunc_snorn_trunc is that the SFH is not limited to being unimodal, and an old population can be fitted with the SFR rising again at recent times.
@@
Table 3: Free and fixed parameters in each SED fitting configuration. Zmin refers to the minimum metallicity available from the SPL used.
@@
The parameters used for fitting in each configuration are outlined in Table 3. The four free dust parameters are applied identically across all implementations, and in addition the fiducial and constant Z implementations have an additional four free SFH parameters and one free metallicity parameters, the $Z_{\odot}$ implementation only has an additional four free SFH parameters, and the b5 SFH implementation has five free SFH parameters and one free metallicity parameter. When fitting the metallicity for each galaxy, we determined that it was necessary to exclude any templates in the fitting with $Z<10^{-4}$. Not only does this ensure a more even comparison between the different SPLs (as some libraries do not extend very low in metallicity space), but also because we found that the inclusion of these low metallicity templates in the SED fitting often resulted in artefacts, made evident by unphysically high peaks in the SFHs at ancient times. This indicates the possibility that the lowest metallicity templates made available by SPLs such as ProGeny and BPASS still suffer from uncertainties and unphysical properties.
The fitting tool Highlander [^5] https://github.com/asgr/Highlander is used for all SED fitting, as done in recent applications of ProSpect by Thorne et al. (2021, 2022a, 2022b, 2023). This tool alternates between the use of the genetic algorithm cmaeshpc [^6] https://github.com/asgr/cmaeshpc and the MCMC (Markov Chain Monte Carlo) CHARM [^7] Component-wise hit and run metropolis algorithm within the LaplacesDemon [^8] https://cran.r-project.org/web/packages/LaplacesDemon/index.html package to first find a solution quickly, and then sample over parameter space to characterise local likelihood space and define sampling uncertainty. Two CMA$+$CHARM runs with 1000 steps each are used to sample parameter space for each galaxy. Rather than implement a fixed burn-in, all steps with a resulting likelihood
@@
are discarded. Typically this discards around 20% of the chain.
@@ 4 Results
We find that the CB19 SPL requires the most computational time (as fitting time scales with the product of the number of age and metallicity bins in the SPL, and CB19 has the greatest number with 221×1522115221$\times$ 15221 × 15999Table 4 from Robotham & Bellstedt (in prep) presents the $N_{\rm Age}$ and $N_{\rm Z}$ numbers for each SPL. ), with M05, and BPASS requiring 0.6 dex less computation time. M05 and BPASS typically require only 3-5 minutes to fit, ProGeny, BC03 and FSPS require 5-10 minutes of fitting time, whereas CB19 requires up to half an hour for a single fit.
@@
The average quality of fit for the SED is shown in Fig. 2 for each SPL, and it can be seen that except for M05, for which the resulting SED fits have a slightly lower likelihood, the quality of the resulting fit is comparable. As discussed by Robotham & Bellstedt (in prep), M05 differs from the other SPLs in that it most notably lacks the UV upturn.
@@
SPL-dependent trends exist in different parts of wavelength space. For example, M05 consistently overestimates the flux in the near-infrared ($Y$, $J$, $H$, and $K$ bands), whereas in this same wavelength range, BC03 tends to slightly underestimate the flux. In the optical bands, most SPLs tend to produce very consistent SEDs however the exception to this is BPASS and ProGeny, which produce notably more flux in the $r$ band than all other SPLs.

### v3

@@ Abstract
We use a volume-complete sample of $\sim$8,000 galaxies from the GAMA survey to characterise the impact of stellar population libraries (SPLs) and model configurations on the resulting inferred galaxy properties from Spectral Energy Distribution (SED) fitting. We compare a fiducial SPL from ProGeny (a new tool that can generate SPLs quickly and flexibly) against five other commonly used SPLs using the SED-fitting code ProSpect. The impact of selecting each SPL is compared to the consequence of changing the model implementation in the SED fitting process, including the implementation of metallicity evolution versus a fixed or constant metallicity, and a functional parametric star formation history (SFH) versus a stepwise parametric (or “non-parametric”) SFH. Furthermore, we use ProGeny to assess the impact of sub-SPL choices, including isochrone selection, stellar spectra selection, and IMF selection. Through a comparison of derived stellar masses, star formation rates, metallicities, ages, and the inferred cosmic star formation history (CSFH), we rank the impact of varying choices. Overall the assumption of a solar metallicity creates the greatest biases, with a substantial impact also caused by the choice of a specific SPL. To recover a CSFH most consistent with observations, we advocate for the use of the fiducial implementation with a skewed Normal functional form for the SFH, and an evolving metallicity, although we note that all studied SPLs underestimate the peak in the CSFH.
@@ 1 Introduction
Stellar population sysnthesis (SPS) modelling uses knowledge about the evolutionary track of stars (isochrones), combined with a stellar initial mass function (IMF, for example Salpeter, 1955; Kroupa, 2002; Chabrier, 2003) and a set of stellar spectra at every evolutionary stage of a star to build up a stellar population generatively (for example, Tinsley & Gunn, 1976; Tinsley, 1978; Bruzual & Charlot, 1993).
SPS models are capable of producing what are known as Simple Stellar Populations (SSPs), which for a given age and metallicity provide a template for the stellar population. A library of such SSPs should cover the entire stellar population, and is referred to as a Stellar Population Library (SPL). These SPLs (of which many flavours exist, including for example Leitherer et al., 1999; Bruzual & Charlot, 2003; Maraston, 2005; Vazdekis et al., 2016) are used as a foundational input to model the observed Spectral Energy Distributions (SEDs) of galaxies via their broadband photometry, in a process named SED fitting. In SED fitting, a stellar population is described in terms of a star formation history (SFH) to account for its age distribution, and a corresponding metallicity history (ZH). This description is used to recreate the stellar emission of the galaxy by building up the appropriate SSPs. An additional dust component is typically modelled, which facilitates the attenuation of the stellar light, as well as the re-emission of this light in the far-infrared through an energy balance process. Much has been written on the topic of SED fitting, and some excellent reviews have been provided by Walcher et al. (2011); Conroy (2013). Today, many different SED fitting codes exist in the literature, including Magphys (da Cunha et al., 2008), BayeSED (Han & Han, 2012), CIGALE (Boquien et al., 2019), Prospector (Leja et al., 2017), BAGPIPES (Carnall et al., 2018), BEAGLE (Chevallard & Charlot, 2016), and ProSpect (Robotham et al., 2020).
An alternative method to extract physical properties of galaxies is full spectral fitting, which is a fitting of SPLs to spectra rather than broadband photometry. Examples of such techniques include pPXF (Cappellari & Emsellem, 2004), Starlight (Cid Fernandes et al., 2005), and Firefly (Wilkinson et al., 2017, which pursues an entirely non-parametric approach to spectral fitting). Despite having similar goals, the difference in wavelength coverage, required model resolution and computational demands are sufficiently large that spectral fitting and SED fitting are generally treated as distinct techniques, with related but separate challenges.
@@
Comparisons of the galaxy properties to have been derived by different SED fitting codes have been made in the past (such as the extensive analysis by Pacifici et al., 2023), and furthermore many studies have analysed the impact of individual elements of SED fitting. This includes Pforr et al. (2012) who presented an early study assessing the impact of assumptions in SED fitting using mocks, Carnall et al. (2019); Leja et al. (2019); Lower et al. (2020); Suess et al. (2022) who studied the impact of SFH parametrisations on stellar masses using Prospector and BAGPIPES, Thorne et al. (2022a) who studied the impact of the inclusion of AGN as a separate component, Haskell et al. (2024) who studied the impact of star formation burstiness, and Jones et al. (2022) who studied the impact of different dust models.
@@
In this work, we aim to characterise the impact of the selected SPL on the final galaxy properties from SED fitting. We compare the SPL impact to that of other modelling choices, including the SFH parametrisation, and the metallicity modelling implementation. Finally, by generating individual SPLs with typical stellar spectra [^2] In the context of ProGeny “stellar spectra” encapsulates both empirical and theoretical spectra of stellar photospheres., isochrones, and IMFs, we further characterise the impact of these sub-SPL choices on the final galaxy properties derived using SED fitting. This sub-SPL comparison is facilitated through the development of a new SPL generation software ProGeny, which is outlined in a companion paper by Robotham & Bellstedt (2024).
@@ 2.1 GAMA
An overview of the physical properties of the selected sample is provided in Fig. 1. The stellar mass is plotted against the redshift, sersic index, and physical size. Stellar masses, sersic indices and sizes have all been derived by Bellstedt et al. (2024) using ProFuse, where 2D light distributions have been simultaneously modelled across optical and near-infrared bands to derive the structural and SED-derived properties self-consistently. Note that the plotted stellar masses are independent from those measured in subsequent sections of this paper.
Figure 1: Properties of the GAMA sample analysed in this work, as derived using the single-sersic mode of ProFuse (Bellstedt et al., 2024). We also present the Vmax corrections that are computed to account for incompleteness in the sample.
Vmax values (which describe the maximum volume to which each galaxy is detectable within the sample selection) for each galaxy in the sample have been computed using ProSpect to identify the maximum redshift where the galaxy falls within the $m_{r}<19.65$ limit. This has been done by directly shifting the redshift of the galaxy spectrum without considering any galaxy evolution (i.e., the assumed SFH remains the same). For galaxies that have a Vmax value less than the volume covered by the sample, a completeness correction can be defined as $V_{\rm survey}/V_{\rm max}$. The corrections computed for individual galaxies are presented in the fourth panel of Fig. 1.
@@ 2.2 Stellar Population Libraries
SPL Reference Isochrone Stellar Spectra IMF Z Range
@@
M05 Maraston (2005)∗∗ Frascati, Geneva 1992 BaSeL, TP-AGB Kroupa (2002) [1e-3, 0.04]
FSPS Conroy et al. (2009) Padova MILES, BaSeL Chabrier (2003) [2e-4, 0.03]
@@
ProGeny Robotham & Bellstedt (2024) MIST C3K Chabrier (2003) [2e-6, 0.063]
  • Isochrones: Padova 1994 (Girardi et al., 2002), Geneva (Lejeune & Schaerer, 2001), Frascati (Cassisi & Salaris, 1997), Geneva 1992 (Schaller et al., 1992), Padova (Marigo & Girardi, 2007; Marigo et al., 2008), Cambridge STARS (Eldridge et al., 2008), PARSEC-COLIBRI (Marigo et al., 2017), MIST (Dotter, 2016).
  • Stellar spectra: BaSeL (Lejeune et al., 1997), STELIB (Le Borgne et al., 2003), Pickles (Pickles, 1998), MILES (Sánchez-Blázquez et al., 2006), C3K (Conroy et al., 2018), TP-AGB (Lançon & Mouhcine, 2002).
  • Comments:
  • Note that the BC03hr version of BC03 was used in this work.
  • ∗∗ The M05 version used includes red horizontal branch stars at all metallicities, resulting in no UV source at old ages.
  • Here the stellar spectra from MILES are used, but the wavelength extension into the UV and NIR is provided by BaSeL.
  • CB19 provides an evolution of BC03, including updated spectra for hot massive stars that produce more He II-ionizing radiation, intended to better reproduce observations (Plat et al., 2019).
The six SPLs compared in this work are presented in Table 1, and include (in order of original release) BC03 (Bruzual & Charlot, 2003), M05 (Maraston, 2005), FSPS (Conroy & Wechsler, 2009), BPASS (Stanway & Eldridge, 2018), CB19 (an update of Bruzual & Charlot 2003, and introduced in Plat et al. 2019), and ProGeny [^4] Available publicly at https://github.com/asgr/ProGeny Robotham & Bellstedt (2024) [^5] While available for use within ProGeny, we have elected not to include the libraries EMILES (Vazdekis et al., 2016) and XSL (Verro et al., 2022) in this paper, due to their truncated wavelength range, and lack of availability of low-age templates. In general, these libraries are better suited for spectral analysis, rather than broadband SEDs. . Each SPL has been constructed using stellar spectra and isochrones that originate from slightly different sources. Furthermore, the nature by which these libraries have been constructed (including interpolation schemes, and how different stellar spectra sets are prioritised) also vary. For further details on the exact construction of these SPLs, please see Robotham & Bellstedt (2024).
@@ 3 Method
We use the SED fitting code ProSpect (Robotham et al., 2020) to test and compare the relative impact of using each of the independent stellar population libraries (SPLs herein), the metallicity implementation (as this is a parameter that varies significantly in terms of its degeneracy with age with different SPLs), and broadly the SFH parametrisation. When fitting the photometric data from GAMA, we augment the measured uncertainties with the error floors presented in table 1 of Bellstedt et al. (2020b), to account for systematic photometric and modelling uncertainties.
@@
The final b5 SFH configuration employs the same metallicity evolution implementation as the fiducial model, however it employs a stepwise parametrisation for the star formation history (available in ProSpect as the massfunc_b5 function), more similar to the stepwise parametric SFHs used typically by codes like Prospector (Leja et al., 2017). This parametriation divides the SFH into five time bins, where the star formation rate (SFR) is fitted independently in each bin. The massfunc_b5 parametrisation is identical to that outlined by Robotham et al. (2020). The five bins are divided by six lookback time boundaries relative to the observed epoch of the galaxy, at 0, 0.1, 1, 5, 9 in Gyr, and the upper boundary of the oldest bin is defined as 13.38 Gyr in absolute lookback time (which shifts relative to the SFH of any individual galaxy observed at different redshifts). These bin limits have been determined so that a constant shift in SFR would have a similar impact on the bolometric light output from each bin, meaning that a modification of SFR in each bin has a roughly similar impact on the resulting SED. The advantage of this over massfunc_snorn_trunc is that the SFH is not limited to being unimodal, and an old population can be fitted with the SFR rising again at recent times.
@@
Table 3: Free and fixed parameters in each SED fitting configuration. Zmin refers to the minimum metallicity available from the SPL used. The Log column indicates whether the parameter was fitted in log space or not, and if no response is indicated then the parameter was set to be fixed (not free).
@@
The parameters used for fitting in each configuration are outlined in Table 3. The four free dust parameters are applied identically across all implementations, and in addition the fiducial and constant Z implementations have an additional four free SFH parameters and one free metallicity parameter. The $Z_{\odot}$ implementation only has an additional four free SFH parameters on top of the free dust parameters, and the b5 SFH implementation has five free SFH parameters and one free metallicity parameter in addition to the free dust parameters. When fitting the metallicity for each galaxy, we excluded any templates in the fitting with $Z<10^{-4}$ to ensure a more even comparison between the different SPLs. Note that in our use of ProSpect, we impose no restrictions on the age/metallicity combinations made available by the chosen SPL.
The optimizer Highlander [^6] https://github.com/asgr/Highlander is used for all SED fitting, as done in recent applications of ProSpect by Thorne et al. (2021, 2022a, 2022b, 2023). This tool alternates between the use of the genetic algorithm cmaeshpc [^7] https://github.com/asgr/cmaeshpc and the MCMC (Markov Chain Monte Carlo) CHARM [^8] Component-wise hit and run metropolis algorithm within the LaplacesDemon [^9] https://cran.r-project.org/web/packages/LaplacesDemon/index.html package to first find a solution quickly, and then sample over parameter space to characterise local likelihood space and define sampling uncertainty. Two CMA$+$CHARM runs with 1000 steps each are used to sample parameter space for each galaxy. Rather than implement a fixed burn-in, all steps with a resulting likelihood
@@
are discarded, where LP is the log posterior of each step. Typically this discards around 20% of the chain.
@@ 4 Results
Examples of SED fitting outputs for a single galaxy (CATAID 7623101010Detailed information on this source is available through the GAMA Single Object Viewer https://www.gama-survey.org/dr4/tools/sov.php. ) are provided in Fig. 2. The fiducial configuration has been used in the upper plot of Fig. 2 for each SPL, whereas the b5 configuration has been presented in the lower plot. These examples provide an indication of the general quality of fit given the data quality, with the typical fitted range in star formation and metallicity histories. Differences in the SFHs are seen when varying the SPLs, but also (unsurprisingly) when changing the SFH parametrisation. It is notable that the metallicity histories are generally less affected by the SFH parametrisation change than the SPL change in this example.
Figure 2: The ProSpect output for galaxy 7623 with the fiducial configuration (top) and b5 configuration (bottom), for each SPL analysed in this work.
@@
We find that the CB19 SPL requires the most computational time (as fitting time scales with the product of the number of age and metallicity bins in the SPL, and CB19 has the greatest number with 221×1522115221$\times$ 15221 × 15111111Table 4 from Robotham & Bellstedt (2024) presents the $N_{\rm Age}$ and $N_{\rm Z}$ numbers for each SPL. ), with M05, and BPASS requiring 0.6 dex less computation time. M05 and BPASS typically require only 3-5 minutes to fit, ProGeny, BC03 and FSPS require 5-10 minutes of fitting time, whereas CB19 requires up to half an hour for a single fit.
@@
The average quality of fit for the SED is shown in Fig. 4 for each SPL. The shaded region is purely for reference, indicating ±$\pm 0.1$ dex (and this shading is used similarly for subsequent plots comparing parameter offsets between SPLs). It can be seen that except for M05, for which the resulting SED fits have a slightly lower likelihood, the quality of the resulting fit is comparable. As discussed by Robotham & Bellstedt (2024), M05 differs from the other SPLs in that it most notably lacks the UV upturn [^12] This is a consequence of the standard M05 version used, which uses red horizontal branch (HB) stars at all metallicities, rather than the hot HB (sometime referred to as extreme HB or eHB) stars at high metallicity, which would produce a UV source for old ages. .
@@
SPL-dependent trends exist in different parts of wavelength space. For example, M05 consistently overestimates the flux in the near-infrared ($Y$, $J$, $H$, and $K$ bands), whereas in this same wavelength range, BC03, CB19, and FSPS tend to slightly underestimate the flux. In the optical bands, most SPLs tend to produce very consistent SEDs however the exception to this is BPASS and ProGeny, which produce notably more flux in the $r$ band than all other SPLs.